Assume the random variable x is normally distributed with mean μ=50 and standard deviation σ=7. Find the indicated probability. P(x>34)
Answer #1

Given : mean = 50, standard deviation = 7

Using normal distribution,

P( x < a) = \phi(\frac{a- mean}{\sigma})

Thus, P( x > 34) = 1 - P( x < 34)

= 1 - phi{ ( 34 - 50)/7}

= 1 - phi(-2.28)

= 1 - 0.0113

= 0.9887

Please like the answer,Thanks!